Allocation: Fixed and Variable Costs, Budgeted Fixed and Variable Costs

Biotechtron, Inc., has two research laboratories in the Southwest, one in Yuma, Arizona, and the other in Bernalillo, New Mexico. The owner of Biotechtron centralized the legal services function in the Yuma office and had both laboratories send any legal questions or issues to the Yuma office. The legal services support center has budgeted fixed costs of $177,000 per year and a budgeted variable rate of $68 per hour of professional time. The normal usage of the legal services center is 2,340 hours per year for the Yuma office and 1,560 hours per year for the Bernalillo office. This corresponds to the expected usage for the coming year.

Required:

1. Determine the amount of legal services support center costs that should be assigned to each office. In your computations, carry ratio values out to two decimal places. Round your final answers to the nearest dollar if rounding is required.

Yuma $
Bernalillo $

2. Since the offices produce services, not tangible products, what purpose is served by allocating the budgeted costs?

It serves the same purposes as costing out tangible products (e.g., pricing, profitability analysis, and performance evaluation).

3. Now, assume that during the year, the legal services center incurred actual fixed costs of $180,400 and actual variable costs of $254,440. It delivered 3,730 hours of professional time—2,520 hours to Yuma and 1,210 hours to Bernalillo.

Determine the amount of the legal services center's costs that should be allocated to each office.

Yuma $
Bernalillo $
Answer #1

1.

Yuma = 2,340*$68+$177,000*2,340 / 3,900 (2,340+1,560)

Yuma = $159,120+(177,000*2,340 / 3,900)

Yuma = $159,120+106,200 = $265,320

Bernalillo =  1,560*$68+$177,000*1,560 / 3,900 (2,340+1,560)

Bernalillo = $106,080+(177,000*1,560 / 3,900)

Bernalillo = $106,080+70,800 = $176,880

3.

Yuma = 2,520*$68+$106,200

Yuma = $171,360+106,200 = $277,560

Bernalillo = 1,210*$68+$70,800

Bernalillo = $82,280+70,800 = $153,080